• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 08:26:02 on July25,2003

    Entry number 106775

    keyword=Owl shift summary

    Taking over with owl shift
    Shift crew :
    Alexandre Camsonne
    Jennifer Niedziela
    Alexandre Deur
    Jing Yuan (momentarily for preshower and shower)

    00:00 Energy pass change


    1:18 Run 22754/2778 ended with 245K events of cosmics data

    Jing Yuan working on preshower switching HV of preshower blocks on and off.

    Run 22755/2779 Junk cosmics run ( HV were changed in the middle )

    Run 22756/2780/3624 Cosmic run with right arm preshower 24,shower 22,36,62 HV off

    2:14 start Run 22757/2781/3625 Cosmics run
    Turn on the right arm HV :preshower #24(-1077);shower#22(-1850),#36(-1600),#62(-1600).
    for shower :100 higher HV than before

    3:15 MCC calls to send some test beam in Hall

    4:50 Beam steered to (0,0)
    Asked for 5 uA raster off for harp scan

    4:20 Beam seems OK but they didn't go through the Compton chicane
    Asked to go through it and will come back to the position tuning

    5:00 Compton tune done
    working on beam profile and position
    beam dispersion : 68 um in X and 6 um in Y

    7:10 Moved to multifoil C target
    Beam width on 03B is (80,55) um best that could be done
    7:15 Start run

    7:30 Raster on

    7:50 Target moved to single foil


    Run summary
    22754/2778 cosmics run 245K events
    22755/2779 junk run
    27756/2780/3624 cosmics preshower 24, shower 23,36,62 off
    27757/2781/3625 cosmics shower and preshower back on
    22758/2782/3626 run taken while beam tuning
    22758/2782/3626 run taken while beam tuning
    22787/2787/3628 acceptance run 3M/140 K multifoil raster off
    22789/2789/3629 acceptance run 3M/150 K multifoil raster on
    22790/2790/3630 acceptance run 2.5M/223K single foil raster off (junk run someone hit a ARC scan during the run)